The table shows cost of purchasing textbooks. What is cost of textbook and cost of shipping when 2 textbooks have a total cost of $190. 4 textbooks would total cost $350.
Y=160 (x)+b is so far what I got.

Answer by Boreal(15235)

Books =x
shipping = y
2x+y=190
4x+y=350
=========
2x+y=190
-4x-y=-350
-2x= -160
x=$80
==========
y=$30
=========